OpenStudy (anonymous):

Solved the system by substitution.if the system is inconsistent and has no solution,state this .if the system is dependent,write the form of the solution for any real number x.3x+12y=10 x+4=7

OpenStudy (anonymous):

\[1)3x+12y=10\] \[2)x+4=7\] \[1)3x+12y=10\] \[2) x = 7-4\] \[1)3x+12y=10\] \[2)x=3\] \[1)3(3)+12y=10\] \[2)x=3\] \[1)9+12y=10\] \[2)x=3\] \[1)12y=1\] \[2)x=3\] ANSWER: \[1)y=1/2\] \[2)x=3\]

OpenStudy (anonymous):

is the solution consistence or has no solution or dependent

OpenStudy (anonymous):

Don't know what the question is really asking, but thats how you solve the equations

OpenStudy (anonymous):

This was the question : Solved the system by substitution.if the system is inconsistent and has no solution,state this .if the system is dependent,write the form of the solution for any real number x

OpenStudy (anonymous):

Then it is dependant, the forms are the 2 answers i gave you

OpenStudy (anonymous):

I dont know after arriving at the final answer did i have to say its inconsisten or has no solution

OpenStudy (anonymous):

THats my best guess

OpenStudy (anonymous):

Are you familiar with if an answer is a solution,Inconsistent or dependent ,Because I have solved some equation but I dont know if they are dependent or independent or inconsistent

OpenStudy (anonymous):

Heh, I havent taken algebra in forever. Look to whatever you teacher told you and look at my answer and try to guess it.

OpenStudy (anonymous):

ok